# Surfel-based Gaussian Inverse Rendering for Fast and Relightable Dynamic Human Reconstruction from Monocular Videos

### v2

@@ Abstract
Efficient and accurate reconstruction of a relightable, dynamic clothed human avatar from a monocular video is crucial for the entertainment industry. This paper introduces the Surfel-based Gaussian Inverse Avatar (SGIA) method, which introduces efficient training and rendering for relightable dynamic human reconstruction. SGIA advances previous Gaussian Avatar methods by comprehensively modeling Physically-Based Rendering (PBR) properties for clothed human avatars, allowing for the manipulation of avatars into novel poses under diverse lighting conditions. Specifically, our approach integrates pre-integration and image-based lighting for fast light calculations that surpass the performance of existing implicit-based techniques. To address challenges related to material lighting disentanglement and accurate geometry reconstruction, we propose an innovative occlusion approximation strategy and a progressive training approach. Extensive experiments demonstrate that SGIA not only achieves highly accurate physical properties but also significantly enhances the realistic relighting of dynamic human avatars, providing a substantial speed advantage. We exhibit more results in our project page: https://GS-IA.github.io.
@@ 1 Introduction
Reconstructing a clothed human avatar and estimating its Physically-Based Rendering (referred to as PBR) properties from monocular videos is a challenging task in computer vision and graphics. This process involves disentangling geometry, PBR materials, and environment illumination. With accurate PBR properties, avatars can be deformed into different poses and re-rendered under novel illuminations, as shown in Fig. 1, offering significant potential for virtual reality, movies, and games.
@@
Early works aimed to estimate PBR properties from single images [1, 2, 3, 4, 5, 6, 7] or multi-view inputs [8, 9, 10, 11, 12, 13, 14, 15], often assuming known lighting environment maps or scene geometry, which is impractical. Radiance fields [16] have emerged as effective 3D-consistent scene representations. Works in this direction [17, 18, 19, 20, 21, 22, 23, 24] estimate geometry, material, and environment light for static scenes from multi-view images but do not account for dynamics, crucial for modeling clothed human avatars. Recent research [25, 26, 27] has advanced clothed human animation by modeling radiance fields and PBR properties for dynamic humans, but NeRF optimizes density rather than geometry directly [28], leading to suboptimal geometry quality. Some works explore neural implicit surfaces [29, 30] for better geometry, but all these methods rely on dense point sampling from MLPs [16] or multi-resolution hash-grids [31], slowing down training and rendering due to extensive query operations.
@@
Recently, Gaussian Splatting [32, 33] has revolutionized high-quality, real-time rendering through its explicit representation and efficient splatting-based rasterization on modern GPUs. Unlike prior methods [34, 35, 29] that depend on complex backward skinning [36, 37] for transformations to a canonical space [38], Gaussian primitives for clothed human avatars are easily and efficiently animated via forward skinning, similar to the template mesh [39]. Recent studies [40, 41, 42, 43, 44, 45, 46] have applied Gaussian splatting techniques to reconstruct dynamic clothed human avatars from monocular or multi-view videos, achieving remarkable rendering quality and speed. However, these approaches are limited to capturing radiance fields and fail to recover the full Physically-Based Rendering (PBR) properties of dynamic humans, thus lacking generalization to novel illuminations.
@@
Our goal is to efficiently and accurately reconstruct a relightable, dynamic clothed human avatar from a monocular video, addressing two key challenges: i.e., disentangling coupled materials and shadow effects caused by lighting occlusion, and achieving high-quality geometry of the clothed human avatar. To this end, we propose the Surfel-based Gaussian Inverse Avatar (SGIA), inspired by the efficiency of Gaussian representation. SGIA recovers both the radiance field and Physically-Based Rendering (PBR) properties of clothed human avatars from monocular videos, enabling avatars to adapt to any scene and varying lighting conditions. We define a set of canonical PBR-awre 2DGS in a canonical space, which are transformed into the observation space via forward linear blend skinning based on different poses. To accurately model the dynamics of clothed human avatars, we combine articulation from the template model Skinned Multi-Person Linear Model (SMPL)[39] with learnable latent bones conditioned on human poses, drawing inspiration from[40].
@@
To effectively disentangle the coupled materials and shadow effects caused by lighting occlusion, inspired by previous works [47, 48, 49, 50], we add learnable parameters, such as albedo, roughness, and metallic, to each Gaussian primitive for the clothed human avatar. While occlusion modeling can be efficiently handled with spherical harmonics in static scenes, it becomes challenging in dynamic scenes where occlusion relationships change over time. Directly calculating occlusion with ray tracing [51] for each Gaussian is accurate but inefficient. To achieve efficient occlusion calculations, we use modern ray-casting algorithms on the limited vertices of the template mesh. Our hybrid strategy combines point and mesh benefits: querying the nearest points on the template mesh for each Gaussian and using surface points and normal vectors to bake ambient occlusion from the template mesh. Given the minimal occlusion difference between the clothed human and the unclothed SMPL [39], our method balances efficiency and accuracy, effectively disentangling materials and lighting.
@@
To accurately reconstruct the surface geometry of clothed human avatars, which is crucial for physically based inverse rendering, we propose to leverage the 2D Gaussian Splatting (referred to as 2DGS) [33] as our primitive representation due to its surface-aligned nature, improving geometry reconstruction quality. Each 2D surfel’s orientation serves as a local normal direction. Initially, we optimized the rough shape using the original Gaussian color. However, the geometry remains inaccurate after image reconstruction, especially under high light conditions (Fig.3). Unlike previous works[47, 50, 49] that freeze Gaussian shape after the first stage, we refine the shape attributes during the PBR optimization stage, leveraging the normal-correlated appearance to enhance geometry. While the original normal consistency loss in 2DGS maintains a splat normal consistency with geometry, it can hinder normal optimization. To address this, we propose a progressive optimization strategy that achieves both accurate geometry and consistent splat normals.
@@
By utilizing the materials, occlusion, and geometry for each splat, we calculate the PBR color of each Gaussian in the observation space and apply the splatting process [52, 33] designed for 2DGS to produce the PBR image. To address the ambiguity between lighting and materials and achieve plausible PBR material solutions, we use regularization terms to maintain material smoothness on planar surfaces and ensure natural lighting. Extensive experiments on synthetic and real-world datasets show that our method enables fast, relightable clothed human avatar reconstruction within 40 minutes and renders at $540\times 540$ resolution at 5 FPS, which is 100×100$\times$100 × faster than state-of-the-art NeRF-based human inverse rendering methods. Quantitative evaluations demonstrate that our method achieves comparable PBR property estimation accuracy to state-of-the-art methods. Comprehensive ablation studies further validate our approach.
@@ 2 Related Work
2.1 Inverse rendering for static scenes.
@@
Estimating the physical properties from images is a highly indeterminate problem due to its under-constrained nature. Previous works on BTF [53, 54, 55, 56, 57] and SVBRDF [58, 11, 14, 59] estimation relay on specific assumptions of specific viewpoints, lighting patterns, or capture setups. Other works explore inverse rendering on single images [1, 2, 3, 4, 5, 6, 7] or multi-view inputs [8, 9, 10, 11, 12, 13, 14, 15]. Recovering shape and illumination from RGB images is heavily ill-posed due to the coupling of shape, materials, and lighting in real-world captured images. These works can only recover high-quality material for each view but fail to achieve realistic and consistent multi-view relighting results under real-world natural illuminations due to the lack of 3D consistent material fields.
@@
With the rapid development of Neural Radiance Fields (referred to as NeRFs) [16], lots of works [23, 17, 18, 19, 21, 20, 60, 61, 22, 62, 63, 64] based on neural field have been proposed to achieve inverse rendering from multi-view images. Specifically, NeRFactor [23] enabled full estimation of the physical properties (geometry, albedo, materials, and lighting) by incorporating the effect of shadows. Neural-PIL [21] explored to approximate pre-integration light integration process by a neural network [65] to achieve fast calculation of outgoing radiance. PhySG [22] proposed to use mixtures of spherical Gaussians to approximate the light transport efficiently. InvRender [62] proposed to model the indirect illumination by distilling a pre-trained NeRF into MLPs to achieve more accurate light disentangling. TensoIR [24] took advantage of tensor decompostion [66] to improve the reconstruction quality. Due to the inaccurate geometry modeling in NeRFs, some other works [67, 68, 69, 70, 71] explored neural signed distance function (SDFs) rendering to achieve more accurate geometry reconstruction during inverse rendering. Among them, NeRO [67] applied the split-sum to constrain the learning of reflective objects, and it can achieve high-quality geometry reconstruction for the high reflective objects. The deep Marching Tetrahedra [72] provided an efficient and differentiable mesh extraction from neural implicit surface representation. With the benefit of differentiable mesh representation and the differentiable mesh rasterizer, Nvdiffrec [73] proposed to conduct light integration on the mesh surface to achieve rapid and photo-realistic physically based rendering. NvdiffrecMC [74] showed that by exploring the combination of mesh-based Mote-Carlo ray-tracing and multi-importance sampling, the decomposition into shape, materials, and lighting can be further improved.
@@
More recently, Gaussian Splatting based methods [32, 75, 33, 76] have achieved high-quality rendering and real-time rendering speed thanks to the forward mapping splatting process. Some works [47, 48, 49, 50] explored to use either the learnable normal vector [47, 50] or the shortest axis of the Gaussian [48, 49] to achieve geometry representation for inverse rendering. However, these works only focused on the inverse rendering of the static scene, while our work focuses on the inverse rendering of a dynamic clothed human avatar, which is more challenging due to the avatar’s dynamic.
2.2 Clothed human avatar modeling.
@@
In addition to the textured mesh, NeRFs [16] also became a useful representation for photo-realistic clothed avatar reconstruction [77, 38, 78, 79, 80, 35, 34, 81, 82, 83, 84] from monocular or multi-view videos. NeuralBody [38] assigned the latent code on the canonical space SMPL [39]vertices and applied the LBS [39] to transform the clothed human avatars to articulation space. For the ray sampling in NeRF [16], the world space points are first transformed to articulated SMPL space to query the radiance. They also rely on time-varied latent to fit the pose-dependent deformation. Animatable NeRF [77] further proposed a pose-dependent deformation network to fit the clothed human conditioned on human. With pose-dependent modeling, the reconstructed avatar can be generalized to novel poses. They also utilize inverse LBS, which can transform the sample points from SMPL articulated space to the canonical space to query the radiance, thus achieving accurate and fast radiance field modeling for humans. However, there exists a large ambiguity in the inverse LBS calculation. InstantAvatar [35] proposed to incorporate the accurate and root finding [37, 36] for inverse LBS and multi-resolution Hash Grids for radiance fields modeling. Recent advances in Gaussian Splatting [32] assigned the radiance to explicit points that can be easily animated by manipulating the skinning weight from template mesh. The forward splatting process does not need the transformation from articulated space to canonical space. Therefore, both the ambiguous inverse LBS and time-consuming root-finding in Snarf can be avoided. Most works [43, 40, 44, 85, 86] explored similar techniques to achieve fast clothed human avatar reconstruction from monocular videos. In contrast with these works, our work aims to model the PBR properties rather than only the radiance. We also take the disentangling of geometry, materials, and environmental lighting to achieve estimation of full PBR properties.
2.3 PBR properties reconstruction of clothed human avatars.
@@
Previous works [87, 88, 89, 90] explored neural networks to learn the decomposition of materials and lighting from single portrait images. Recently, High quality physical properties reconstruction of clothed human avatar [30, 91, 25, 26, 29, 27, 92, 93, 94, 79] have been widely explored with the development of Neural Radiance Fields [16]. DS-NeRF [79] proposed to query the lighting at world space to disentangle the light and texture. Relighing4D [25] first explored the MLPs representation based on NeuralBody [38] to jointly estimate the shape, lighting, and the albedo of dynamic humans from monocular or sparse videos under unknown illuminations by disentangling visibility via MLPs. RANA [92] pre-trained a mesh representation on ground truth 3d assets to approximate the radiance transfer under different lighting. Relightable Avatar [26] pretrained an occlusion approximation neural network from a large-scale motion sequence dataset [95, 96]. Intrinsic Avatar [29] implemented a fast secondary ray-tracing at articulated space to calculate the visibility and the indirect illuminations. Different from these works, our method explores a relightable clothed human avatar representation based on Gaussian splatting [32] to achieve fast training and fast rendering. Concurrent with our work, Animatable Gaussians [97] proposed a relighting extension that predicts occlusion based on the human pose and position map with the 2D StyleUnet [98] to achieve high-quality rendering. However, their work relied on multi-view videos as input and complex poses during training to achieve pose generalization, which is not suitable for our setting. PhysAvatar [99] also explored the relightable Gaussian Avatar, but it relied on the multi-view videos as input to reconstruct the mesh at the first timestamp.

### v3

@@ Abstract
Efficient and accurate reconstruction of a relightable, dynamic clothed human avatar from a monocular video is crucial for the entertainment industry. This paper presents SGIA (Surfel-based Gaussian Inverse Avatar), which introduces efficient training and rendering for relightable dynamic human reconstruction. SGIA advances previous Gaussian Avatar methods by comprehensively modeling Physically-Based Rendering (PBR) properties for clothed human avatars, allowing for the manipulation of avatars into novel poses under diverse lighting conditions. Specifically, our approach integrates pre-integration and image-based lighting for fast light calculations that surpass the performance of existing implicit-based techniques. To address challenges related to material lighting disentanglement and accurate geometry reconstruction, we propose an innovative occlusion approximation strategy and a progressive training approach. Extensive experiments demonstrate that SGIA not only achieves highly accurate physical properties but also significantly enhances the realistic relighting of dynamic human avatars, providing a substantial speed advantage. We exhibit more results in our project page: https://GS-IA.github.io.
@@ 1 Introduction
Reconstructing a clothed human avatar and estimating its Physically-Based Rendering (referred to as PBR) properties simultaneously from monocular videos is a challenging task in computer vision and graphics. It involves disentangling geometry, PBR materials, and environment illumination. With accurate PBR properties, avatars can be deformed into different poses and re-rendered under novel illuminations, as shown in Fig. 1, offering significant potential for virtual reality, movies, and games.
@@
Early works aim to estimate PBR properties from single images [1, 2, 3, 4, 5, 6, 7] or multi-view inputs [8, 9, 10, 11, 12, 13, 14, 15], often assuming known lighting environment maps or scene geometry, which is impractical. Radiance fields [16] have emerged as effective 3D-consistent scene representations. Works in this direction [17, 18, 19, 20, 21, 22, 23, 24] estimate geometry, material, and environment light for static scenes from multi-view images but do not account for dynamics, crucial for modeling clothed human avatars. Recent works [25, 26, 27] have advanced clothed human animation by modeling radiance fields and PBR properties for dynamic humans, but NeRF (Nerural Radiance Fields) optimizes density rather than geometry directly [28], leading to suboptimal geometry quality. Some works propose neural implicit surfaces [29, 30] for better geometry, but all these methods rely on dense point sampling from MLPs [16] or multi-resolution hash-grids [31], slowing down training and rendering due to extensive query operations.
@@
Recently, Gaussian Splatting [32, 33] has revolutionized high-quality, real-time rendering through its explicit representation and efficient splatting-based rasterization on modern GPUs. Unlike prior methods [34, 35, 29] that rely on complex backward skinning [36, 37] for transformations to a canonical space, clothed human avatars based on Gaussian primitives can be animated easily and efficiently through forward skinning, akin to the template mesh [38]. Recent studies [39, 40, 41, 42, 43, 44, 45] have applied Gaussian splatting techniques to reconstruct dynamic clothed human avatars from monocular or multi-view videos, achieving remarkable rendering quality and speed. However, these approaches are limited to capturing radiance fields and fail to recover the full Physically-Based Rendering (PBR) properties of dynamic humans, thus lacking generalization to novel illuminations.
@@
Our goal is to efficiently and accurately reconstruct a relightable and dynamic clothed human avatar from a monocular video, addressing two key challenges: i.e., disentangling coupled materials and shadow effects caused by lighting occlusion, and achieving high-quality geometry of the clothed human avatar. To this end, we propose the Surfel-based Gaussian Inverse Avatar (SGIA), inspired by the efficiency of Gaussian representation. SGIA recovers the radiance field and PBR properties of clothed avatars from monocular videos, enabling adaptation to unknown lighting. We define a set of canonical PBR-aware 2DGS (2D Gaussian Splatting) [33] in canonical space, transformed to observation space via forward linear blend skinning based on poses. To accurately model the dynamics of clothed human avatars, we combine articulation from the template model Skinned Multi-Person Linear Model (SMPL)[38] with learnable latent bones conditioned on human poses, inspired by GART [39].
@@
To effectively disentangle the coupled materials and shadow effects caused by lighting occlusion, inspired by previous works [46, 47, 48, 49], we add learnable parameters, such as albedo, roughness, and metallic, to each Gaussian for the human avatar. Occlusion modeling can be efficiently handled with spherical harmonics in static scenes [47, 46], where the occlusion relationships remain unchanged over time. However, it becomes challenging in dynamic scenes where occlusion relationships change over time. Directly calculating occlusion with ray tracing [50] for each Gaussian is accurate but inefficient. To achieve efficient occlusion calculations, we use modern ray-casting algorithms [51] on the limited vertices of the template mesh. Our hybrid strategy integrates the advantages of both point and mesh representations: we query the nearest points on the template mesh for each Gaussian and utilize surface points along with normal vectors to bake ambient occlusion from the template mesh. Due to the very small occlusion difference between the clothed human and the unclothed SMPL [38], our approach strikes a balance between efficiency and accuracy, effectively disentangling materials from lighting.
@@
To accurately reconstruct the surface geometry of clothed human avatars, which is crucial for physically based inverse rendering, we propose to leverage the 2DGS [33] as our primitive representation due to its surface-aligned nature, improving geometry reconstruction quality. Each 2D surfel’s orientation serves as a local normal direction. Initially, we optimize the rough shape using the original Gaussian color. The geometry remains inaccurate after image reconstruction, especially under high light conditions (Fig.3). Unlike previous works[46, 49, 48] that freeze Gaussian shape after the first stage, we refine the shape attributes during the PBR optimization stage, leveraging the normal-correlated appearance to enhance geometry. While the original normal consistency loss in 2DGS maintains a splat normal consistency with geometry, it can hinder normal optimization. To address this, we propose a progressive optimization strategy that achieves both accurate geometry and consistent splat normals.
@@
By utilizing the materials, occlusion, and geometry for each splat, we calculate the PBR color of each Gaussian in the observation space and apply the splatting process [52, 33] designed for 2DGS to produce the PBR image. To address the ambiguity between lighting and materials and achieve plausible PBR material solutions, we use regularization terms to maintain material smoothness on planar surfaces and ensure natural lighting. Extensive experiments on synthetic and real-world datasets show that our method enables fast, relightable clothed human avatar reconstruction within 40 minutes and renders at $540\times 540$ resolution at 5 FPS, which is $100\times$ faster than state-of-the-art NeRF-based human inverse rendering methods. Quantitative evaluations demonstrate that our method achieves comparable PBR properties estimation accuracy to state-of-the-art methods. Comprehensive ablation studies further validate our approach.
@@ 2 Related Work
2.1 Inverse rendering for static scenes
@@
Estimating the physical properties from images is a highly ill-conditioned problem due to its under-constrained nature. Previous works on BTF (Bidirectional Texture Function) [53, 54, 55, 56, 57] and SVBRDF (Spatially Varying Bidirectional Reflectance Distribution Function) [58, 11, 14, 59] estimation rely on specific assumptions of viewpoints, lighting patterns, or capture setups. Other works explored inverse rendering on single image [1, 2, 3, 4, 5, 6, 7] or multi-view inputs [8, 9, 10, 11, 12, 13, 14, 15]. Recent works [60, 61] have explored the radiance transfer technique [60] or utilized data priors from diffusion model [61] to achieve inverse rendering. These works can only recover high-quality material for each view but fail to achieve realistic and consistent multi-view relighting results under real-world natural illuminations due to the lack of 3D neural fields.
@@
With the rapid development of Neural Radiance Fields [16], lots of works [23, 17, 18, 19, 21, 20, 62, 63, 22, 64, 65, 66] based on neural field have been proposed to achieve inverse rendering from multi-view images. Specifically, NeRFactor [23] enabled full estimation of the physical properties (geometry, albedo, materials, and lighting) by incorporating the effect of shadows. Neural-PIL [21] proposed to approximate pre-integration light integration process with a neural network [67] to achieve fast calculation of outgoing radiance. PhySG [22] proposed to use mixtures of spherical Gaussians to approximate the light transport efficiently. InvRender [64] proposed to derive the indirect illumination from a pre-trained NeRF to achieve more accurate light disentangling. TensoIR [24] took advantage of tensor decomposition [68] to improve the quality of reconstruction. Due to inaccurate geometry modeling in NeRFs, some other work [69, 70, 71, 72, 73] explored neural signed distance function (SDF) rendering to achieve more accurate geometry reconstruction for inverse rendering. Among them, NeRO [69] applied the split-sum to constrain the learning of reflective objects. It achieved high-quality geometry reconstruction for the high reflective objects. The Deep Marching Tetrahedra [74] provided an efficient and differentiable mesh extraction from neural implicit surface representation. With the benefits of differentiable mesh representation and the differentiable mesh rasterizer, Nvdiffrec [75] proposed to conduct light integration on the mesh surface to achieve rapid and photo-realistic physically based rendering. NvdiffrecMC [76] showed that by exploring the combination of mesh-based Mote-Carlo ray-tracing and multi-importance sampling, the decomposition into shape, materials, and lighting can be further improved.
@@
More recently, Gaussian Splatting based methods [32, 77, 33, 78] have achieved high-quality rendering and real-time rendering speed thanks to the forward mapping splatting process. Some works [46, 47, 48, 49] proposed to use either the learnable normal vector [46, 49] or the shortest axis of the Gaussian [47, 48] to achieve geometry representation for inverse rendering. However, these works only focused on the inverse rendering of the static scene, while our work focused on the more challenging inverse rendering of a dynamic clothed human avatar.
2.2 Clothed human avatar modeling
@@
In addition to the textured mesh, NeRFs [16] also became a useful representation for photo-realistic clothed avatar reconstruction [79, 80, 81, 82, 83, 35, 34, 84, 85, 86, 87] from monocular or multi-view videos. NeuS2 [88] achieved high quality human reconstruction from multi-view videos without relying on the human template [38]. NeuralBody [80] assigned the latent code on the canonical space SMPL [38] vertices and applied the LBS [38] to transform the clothed human avatars to articulation space. For the ray sampling in NeRF [16], the world space points are first transformed to articulated SMPL space to query the radiance. They also relied on time-varied latent to fit the pose-dependent deformation. Animatable NeRF [79] further proposed a pose-dependent deformation network to fit the clothed human conditioned on human. They also utilized the inverse LBS [79], which can transform the sample points from SMPL articulated space to the canonical space to query the radiance, thus achieving accurate and fast radiance field modeling for humans. However, there exists a large ambiguity in the inverse LBS calculation. InstantAvatar [35] proposed to incorporate root finding algorithms [37, 36] and multiresolution hash grids to reduce ambiguity in the calculation process. Recent advances in Gaussian Splatting [32] assigned the radiance to explicit points that can be easily animated by manipulating the skinning weight from template mesh. The forward splatting process does not need the transformation from articulated space to canonical space. Therefore, both ambiguous inverse LBS and time-consuming root-finding in Snarf [37] can be avoided. Most works [42, 39, 43, 89, 90, 91] explored similar techniques to achieve fast clothed human avatar reconstruction from monocular videos. Unlike these works, our work aims to model the PBR properties rather than the radiance only. We also consider the disentanglement of geometry, materials, and environmental lighting to achieve estimation of the full PBR properties.
2.3 PBR properties reconstruction of clothed human avatars
@@
Previous works [92, 93, 94, 95] explored neural networks to learn the decomposition of materials and lighting from a single portrait image. Recently, high quality physical properties reconstruction of clothed human avatar [30, 96, 25, 26, 29, 27, 97, 98, 99, 82] have been widely explored with the development of Neural Radiance Fields [16]. DS-NeRF [82] proposed to query the lighting at world space to disentangle the light and texture. Relighing4D [25] is the first to explore the MLPs representation based on NeuralBody [80] to jointly estimate the shape, lighting, and the albedo of dynamic humans from monocular or sparse videos under unknown illumination by disentangling visibility via MLPs. RANA [97] pre-trained a mesh representation on ground truth 3d assets to approximate the radiance transfer under different lighting. Relightable Avatar [26] pretrained an occlusion approximation neural network on a large-scale motion sequence dataset [100, 101]. Intrinsic Avatar [29] implemented a fast secondary ray-tracing at articulated space to calculate the visibility and the indirect illuminations. Different from these works, our method explores a relightable clothed human avatar representation based on Gaussian splatting [32] to achieve fast training and fast rendering. Concurrent with our work, Animatable Gaussians [102] proposed a relighting extension that predicts occlusion based on the human pose and position map with the 2D StyleUnet [103] to achieve high-quality rendering. However, their work relied on multi-view videos as input and complex poses during training to achieve pose generalization, which is not suitable for our setting. PhysAvatar [104] also explored the relightable Gaussian Avatar, but it relied on multi-view videos as input to reconstruct the mesh at the first timestamp.